Could someone help with this problem. Detwmine the equation of the line that has a slope of 5 and passes through the point (1,3). Write the equation in slope intercept form.
Answer by venugopalramana(3286)

see example below and try.if still in difficulty please come back
how do I write an equation in slope-intercept form that satisfies the condition.
slope= -0.5, passes through (2,-3)
std.eqn is
y=mx+c.....m=slope=-0.5
c= y intercept
y= -0.5x +c...it passes through 2,-3..hence
-3= -0.5*2+c= -1+c
c=1-3 = -2
eqn.is
y= -0.5x -2